### v2

@@ 4.3 Discussion of results
$\lambda^{(1)}$ $0.16(-0.15,0.32)$ $0.14\boldsymbol{(-0.18,0.37)}$ $0.15\boldsymbol{(-0.25,0.42)}$
$\lambda^{(2)}$ $0.79(0.57,1.17)$ $0.8\boldsymbol{(0.44,1.24)}$ $0.78\boldsymbol{(0.45,1.23)}$
$\sigma^{(1)}$ $0.93(0.82,1.1)$ $1.01\boldsymbol{(0.77,1.15)}$ $1\boldsymbol{(0.78,1.18)}$
$\sigma^{(2)}$ $0.74(0.49,1.03)$ $0.8\boldsymbol{(0.48,1.09)}$ $0.86\boldsymbol{(0.43,1.14)}$
$M^{(1)}$ $0.26(-0.14,0.44)$ $0.35\boldsymbol{(-0.19,0.58)}$ $0.35\boldsymbol{(-0.22,0.64)}$
$M^{(2)}$ $0.91(0.78,1.25)$ $0.96\boldsymbol{(0.71,1.36)}$ $0.96\boldsymbol{(0.61,1.33)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.63(-1.15,-0.38)$ $-0.66\boldsymbol{(-1.29,-0.24)}$ $-0.63\boldsymbol{(-1.27,-0.31)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.26(0.89,1.92)$ $1.27\boldsymbol{(0.89,2.06)}$ $1.16\boldsymbol{(0.78,2.37)}$
@@ 5 Conclusions and Future Developments
Integrating multiple observational studies to make unconfounded causal or descriptive comparisons of group potential outcomes in large natural populations presents significant challenges, because of the complexities involved in data heterogeneity, selection bias, and the need for accurate balancing across datasets. Recently, Guha and Li (2024) introduced a unified weighting framework designed to address these challenges by extending inverse probability weighting techniques for integrative analyses. To translate this theoretical framework into practice, we have developed the R package WMAP. This software tool is specifically designed for the integrative analysis of user-specified datasets and implements three advanced weighting approaches, i.e., IC (Integrative Calibration), IGO (Integrative Generalized Optimization), and FLEXOR (Flexible Optimization of Weights). These methods enhance the capacity to estimate multiple propensity scores, compute balancing weights for subjects, evaluate effective sample sizes, and derive various estimands of counterfactual group outcomes. The package also includes functionality for calculating bootstrap variability estimates, which are essential for quantifying the uncertainty of the results. In a practical application, we have used WMAP to analyze differences in mean and standard deviation of gene expression levels between two major breast cancer subtypes, invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC). This analysis yielded biologically meaningful results, highlighting the package’s potential for contributing to the understanding of complex biological systems. WMAP’s utility is not limited to observational studies. It is equally applicable to multi-arm randomized controlled trials (RCTs), particularly in scenarios where within-study allocation mechanisms are known. Furthermore, future updates to WMAP will extend its capabilities to support hybrid study designs that combine data from RCTs and retrospective cohorts. Current limitations, such as challenges in handling high-dimensional biomarker data, are actively being addressed in ongoing development. These enhancements will make WMAP an even more powerful tool for advancing integrative causal inference and descriptive analyses across a wide range of scientific and clinical research domains.
@@ Acknowledgments
This work was supported by the National Science Foundation and National Institutes of Health under award DMS-1854003 to SG, award CA249096 to YL, and awards CA269398 and CA209414 to SG and YL.

### v3

@@ 3.3 Discussion of results
$\lambda^{(1)}$ $0.16(-0.1,0.29)$ $0.12\boldsymbol{(-0.2,0.33)}$ $0.15\boldsymbol{(-0.25,0.38)}$
$\lambda^{(2)}$ $0.79(0.55,1.14)$ $0.84\boldsymbol{(0.5,1.23)}$ $0.88\boldsymbol{(0.52,1.24)}$
$\sigma^{(1)}$ $0.93(0.8,1.1)$ $1.02\boldsymbol{(0.77,1.15)}$ $0.98\boldsymbol{(0.75,1.19)}$
$\sigma^{(2)}$ $0.74(0.5,1.07)$ $0.75(0.48,1.05)$ $0.78\boldsymbol{(0.44,1.06)}$
$M^{(1)}$ $0.26(-0.09,0.35)$ $0.32\boldsymbol{(-0.22,0.55)}$ $0.32\boldsymbol{(-0.23,0.5)}$
$M^{(2)}$ $0.91(0.78,1.23)$ $0.96\boldsymbol{(0.57,1.36)}$ $0.96\boldsymbol{(0.65,1.36)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.63(-1.12,-0.38)$ $-0.72\boldsymbol{(-1.27,-0.28)}$ $-0.74\boldsymbol{(-1.3,-0.33)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.26(0.86,1.96)$ $1.36\boldsymbol{(0.92,2.09)}$ $1.25\boldsymbol{(0.87,2.22)}$
4 Simulations
We used the WMAP package to generate simulated datasets for evaluating various weighting strategies used in inferring population-level characteristics across two subject groups. Specifically, we analyzed these datasets using the causal.estimate() function. We briefly describe the generation strategy below and present the results. Following Guha and Li (2024) and in alignment with the motivating TCGA breast cancer studies, we simulated $R=250$ independent datasets, each containing $J=7$ observational studies, $K=2$ groups, and $L=1$ (i.e., univariate) outcome for $\tilde{N}=500$ subjects. The covariate vectors were sampled with replacement from the demo dataset included in the WMAP package, which mimics the structure of the $N=450$ TCGA breast cancer dataset used in the motivating example. To demonstrate the practical application of the WMAP package, we present results from a representative simulation scenario designed to reflect key features of real-world retrospective studies with confounding. This scenario effectively demonstrates the comparative performance of different weighting strategies implemented in the package. For the reported results, we used the procedure in Section 2 to meta-analyze the seven studies within each artificial dataset.
As an initial step for all $250$ artificial datasets, we conducted k-means clustering on the covariates, $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$, from the WMAP demo dataset to identify lower-dimensional structure, grouping them into $Q=12$ clusters with centers $\mathbf{q}_{1},\ldots,\mathbf{q}_{Q}\in\mathcal{R}^{p}$ and allocated covariate counts $m_{1},\ldots,m_{Q}$. For each artificial dataset $r=1,\ldots,250$, containing $\tilde{N}$ patients, we then generated the data as follows:
  1. 1.
    Generate the group proportions in the natural population by drawing $\boldsymbol{\theta}^{(r)}\sim\mathcal{D}_{K}(\mathbf{1}_{K})$, for $K=2$ groups. Define the group-covariate relationships: $\delta^{(r)}{z}(\mathbf{x})\propto 1$ if $z=1$ and $\delta^{(r)}{z}(\mathbf{x})\propto\exp\bigl{(}\omega_{0}^{(r)}+\omega_{1}^{(r)}\sum_{t=1}^{p}x_{t}/\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}\sum_{t=1}^{p}x_{irt}^{(0)}\bigr{)}$ if $z=2$. Here, we set $\omega_{1}^{(r)}=1$ and choose $\omega_{0}^{(r)}$ such that the population average of $\delta^{(r)}_{z}(\mathbf{x}_{ir}^{(0)})$ equals $\theta_{z}^{(r)}$.
  2. 2.
    Covariates  For each subject $i=1,\ldots,\tilde{N}$, select their covariate vector $\tilde{\mathbf{x}}_{i}^{(r)}=(\tilde{x}_{i1}^{(r)},\ldots,\tilde{x}_{ip}^{(r)})^{\prime}$ by sampling with replacement from the $N=450$ demo dataset covariate vectors.
  3. 3.
    1. (a)
      MPS  Define the group-specific study propensities as follows:
      $\log\bigl{(}\delta_{S=s\mid Z=z}(\mathbf{x})/\delta_{S=1\mid Z=z}(\mathbf{x})\bigr{)}=sz\omega_{1}^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)}/\frac{1}{\tilde{N}}\sum_{i^{\prime}=1}^{\tilde{N}}\sum_{t=1}^{p}\tilde{x}_{i^{\prime}t}^{(r)}$
      for $s=2,\ldots,J$ and $z=1,2$. Assuming the group propensity scores are the same as in the natural population, the marginal propensity score (MPS) is given by $\delta_{sz}(\mathbf{x})=\delta_{s\mid z}(\mathbf{x})\delta_{z}(\mathbf{x})$. For each patient $i=1,\ldots,\tilde{N}$, evaluate their probability vector $\boldsymbol{\delta}^{(r)}(\mathbf{x}_{i})=\bigl{(}\delta_{11}^{(r)}(\mathbf{x}_{i}),\ldots,\delta_{JK}^{(r)}(\mathbf{x}_{i})\bigr{)}$.
    2. (b)
  4. 4.
    Subject-specific observed outcomes  Generate
    $Y_{i}^{(r)}\,\Big{|}\,Z_{i}=z_{i}^{(r)}\stackrel{{\scriptstyle\text{indep}}}{{\sim}}\left(z_{i}^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)}+50,\,\tau_{r}^{2}\right),$
Next, we set aside knowledge of the simulation parameters and analyzed each artificial dataset using the procedure described in Section 2 for the IC, IGO, and FLEXOR pseudo-populations, as implemented in the WMAP package. Define percent ESS as the effective sample size (ESS) scaled for 100 participants. For the 250 simulated datasets, the first panel of Figure 1 displays boxplots of the percent ESS for the FLEXOR, IGO, and IC pseudo-populations. The IC and IGO pseudo-populations showed comparable ESS. The FLEXOR pseudo-population, however, consistently achieved substantially higher ESS across all datasets.
We employed the Stage 2 strategy described in Section 2, which is implemented in the WMAP package, to estimate the average treatment effect (ATE), defined as the difference in group-level counterfactual means, $\lambda^{(1)}-\lambda^{(2)}$. This was achieved by making weighted inferences under each method’s pseudo-population, where covariates were balanced to support causal interpretation. Because each estimator targets the ATE within its respective pseudo-population, we evaluated accuracy by comparing the estimated values to their corresponding true ATEs obtained via Monte Carlo simulation. The second and third panels of Figure 1 show the absolute biases and standard deviations of the FLEXOR, IGO, and IC estimators across 250 synthetic datasets. For each dataset and method, these performance metrics were computed using 200 independent bootstrap samples.
Overall, while the IGO and IC weighting strategies showed comparable performance in estimating the average treatment effect (ATE) across the simulated datasets, FLEXOR consistently outperformed both, as shown in Figure 1. Additionally, it achieved lower absolute bias and standard deviation than the competing methods in all 250 artificial datasets. Notably, although IGO weights are theoretically optimal for estimating the ATE under certain idealized conditions, such as homoscedastic outcomes and correct model specification (see Li and Li, 2019, for single studies), these assumptions did not hold in our simulation setting. The strong empirical performance of FLEXOR in this context underscores its robustness and practical value, particularly when theoretical optimality conditions are violated. By placing greater emphasis on the stability and practical feasibility of the estimated balancing weights, FLEXOR offers a robust and reliable approach to ATE estimation in meta-analytical contexts, especially in scenarios where the assumptions underlying asymptotic efficiency may be less tenable.
Figure 1: Boxplots summarizing the results across 250 simulated datasets ($\tilde{N}=500$ subjects each): (i) percentage effective sample size (ESS), (ii) absolute bias, and (iii) standard deviation of the WATE (Weighted Average Treatment Effect) under three pseudo-populations.
@@ 5 Conclusions and Future Developments
Integrating multiple observational studies to make unconfounded causal or descriptive comparisons of group potential outcomes in large natural populations presents significant challenges, because of the complexities involved in data heterogeneity, selection bias, and the need for accurate balancing across datasets. Recently, Guha and Li (2024) introduced a unified weighting framework designed to address these challenges by extending inverse probability weighting techniques for integrative analyses. To translate this theoretical framework into practice, we have developed the R package WMAP. This software tool is specifically designed for the integrative analysis of user-specified datasets and implements three advanced weighting approaches, i.e., IC (Integrative Calibration), IGO (Integrative Generalized Optimization), and FLEXOR (Flexible Optimization of Weights). These methods enhance the capacity to estimate multiple propensity scores, compute balancing weights for subjects, evaluate effective sample sizes, and derive various estimands of counterfactual group outcomes. The package also includes functionality for calculating bootstrap variability estimates, which are essential for quantifying the uncertainty of the results.
In our illustrative application, we used WMAP to analyze simulated gene expression data that mimic differences between two major breast cancer subtypes: invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC). Although based on synthetic data, the analysis yielded patterns consistent with known biological characteristics of these subtypes, demonstrating WMAP’s potential to inform scientific insights in causal genomics. The package supports both observational studies and multi-arm randomized controlled trials (RCTs), especially when within-study treatment assignment mechanisms are known. Looking ahead, future updates will extend WMAP’s functionality to accommodate hybrid study designs that integrate data from RCTs and retrospective cohorts. Current limitations, such as challenges in handling high-dimensional biomarker data, are being actively addressed. These enhancements aim to make WMAP an even more versatile tool for integrative causal inference and descriptive analysis across diverse biomedical research settings.
@@ Acknowledgments
This work was supported by the National Institutes of Health under award DMS-1854003 to SG, award CA249096 to YL, and awards CA269398 and CA209414 to SG and YL.